Operation Summary for November 30, 2009 - December 2, 2009 |
Monday Nov. 30 The day shift began with Operations monitoring the Tevatron (TeV) store 7407 (220.23E30) and the Recycler stash (56.0E10), with the Antiproton Source (Pbar) stacking (23.2E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day shift, Operations reported a MECAR failure at 2:20 PM, which required a reboot. During the evening shift, Operations reported that store 7407 ended in a quench (believed to be due to a cryostat vacuum problem) at 4:11 PM. Operators prepared D0 for an access, which lasted from 5:53 to 6:37 PM. Operators began the TeV turn-on and then put the TeV into HEP shot setup at 11:17 PM. Tuesday Dec. 1 The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (87.8E10), with Pbar stacking (37.5E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established store 7411 at 12:33 AM with an Average Initial Luminosity (AIL) of 259.6E30. During the day shift, the H- Source had vacuum problems, which reduced intensity. TeV personnel conducted end-of-store studies from 1:55 PM to the end of the day shift. During the evening shift, the TeV end-of-store study ended at 6:59 PM. Operations terminated store 7411 at 7:16 PM and then put the TeV into shot setup. Operations established store 7412 at 8:44 PM with an AIL of 179E30. The Meson MTest experiment T992 began taking beam on the evening shift; it was turned off at midnight. Wednesday Dec. 2 The midnight shift began with Operations monitoring store 7412 (107.71E30) and the stash (86.5E10), with Pbar stacking (19.5E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 7412 at 5:02 AM. Operations put the TeV into shot setup and established store 7413 at 6:28 AM with an AIL of 165.7E30. Operators turned off beam for MI, Booster, NuMI, and MiniBooNE and prepared these areas for accesses. The Plans for Wednesday The plans for today are to monitor the store and to conduct necessary repairs in Linac, Booster, MI, NuMI, and MiniBooNE through to later this afternoon and then resume stacking and stashing. |
|
More Information Comments and Suggestions |
| last modified 12/02/2009 email Fermilab |
| Security, Privacy, Legal | |